OpenStudy (anonymous):

I need to find the mean, mode, median, variance and standard deviation of this data: 101,102,100,110,109,109,108,109

OpenStudy (johnweldon1993):

Alright...so lets handle the mode and the median first.. mode = number that appears most often.. median = number in the middle when all numbers are aligned least to greatest... what are those 2?

OpenStudy (anonymous):

mode=109?

OpenStudy (johnweldon1993):

correct :)

OpenStudy (anonymous):

median= 108? or 109?

OpenStudy (johnweldon1993):

Well lets see... From least to greatest we have \[\large 100, 101, 102, 108, 109, 109, 109, 110\] So the middle would be \[\large \cancel{100}, \cancel{101}, \cancel{102}, \color \red{108, 109}, \cancel{109}, \cancel{109},\cancel{ 110}\] When we end up with 2 numbers...what we do ..is average them...add them together and divide by 2...what do we get?

OpenStudy (anonymous):

108.5? :)

OpenStudy (johnweldon1993):

Perfect... Alright...now...we want the mean...do you know how to find that?

OpenStudy (anonymous):

THANKS, no :(

OpenStudy (johnweldon1993):

That's okay hun, that's what we're going to figure out :) So the mean...you know how we just averaged those 2 numbers to find the median? Well the mean...is the average of all your numbers... So what you do...is add up every number you have....and then divide the answer you get by the number of numbers you added... So here...add up all of them...and then divide by 8 since there are 8 numbers...what do we get?

OpenStudy (anonymous):

118.5?

OpenStudy (johnweldon1993):

Hmm...I get something different but lets check I may have made a mistake... \[\large \frac{100 + 101 + 102 + 108 + 109 + 109 + 109 + 110}{8} = \frac{848}{8} = 106\] Look good? :)

OpenStudy (anonymous):

you are so nice :) thankyou for your patience!

OpenStudy (anonymous):

what about the variance :)?

OpenStudy (johnweldon1993):

mmhmm that's the next one...alright...so we have the mean right? For the variance...we need to see how far away EACH number is from that mean... So mean = 106 how far is 100 from 106? well it is -6 right? continue on from there finding all the other numbers so how far is: 100 from 106 101 from 106? 102 from 106? 108 from 106? 109 from 106? 109 from 106? 109 from 106? 110 from 106?

OpenStudy (anonymous):

-5,-4,2,3,3,3,4? :):)

OpenStudy (johnweldon1993):

Great....so now...we take all those differences ...and square them...(raise the to the 2nd power) and then add all the results so we have \[\large (-6^2) + (-5^2) + (-4^2) + (2^2) + (3^2) + (3^2) +(3^2) + (4^2) \] what do we get when we add all those?

OpenStudy (anonymous):

92 right?

OpenStudy (anonymous):

Im sorry 124!!

OpenStudy (johnweldon1993):

sorry hun I was away...umm lets see \[\large (-6^2) + (-5^2) + (-4^2) + (2^2) + (3^2) + (3^2) +(3^2) + (4^2)\] \[\large 36 + 25 + 16 + 4 + 9 + 9 + 9 + 16 = 241\] Is what I get...hmm where did you go off?

OpenStudy (anonymous):

I dont know, I think i am kind of tired, ive been doing math whole day. Sorry haha

OpenStudy (anonymous):

but thank you so much for your patience! how can I find the standard deviation?

OpenStudy (johnweldon1993):

We aren't done with the variance :P Alright...what we do now...with the 214 we have...is divide it by 8 ...(like we did with the mean)....so the variance will be \[\large \frac{241}{8} = ?\]

OpenStudy (anonymous):

30.12 :)?

OpenStudy (johnweldon1993):

ehh we'll keep the 5 at the end for now :P \[\large Variance = 30.125\] so ready for the standard deviation? its soooo difficult......(sarcasm) :P

OpenStudy (anonymous):

HAHAHAH dont make fun of me hahaha you are the worst

OpenStudy (anonymous):

ok please help me do my extra hard algebra so I can finally take a break

OpenStudy (johnweldon1993):

I would never hun :P haha but alright...the standard deviation is just the square root of the Variance \[\large \sqrt{30.125} = ?\]

OpenStudy (anonymous):

5.48?

OpenStudy (johnweldon1993):

we'll round up to 5.49...but yes....And you are all done :)

OpenStudy (anonymous):

THANK YOU SO MUCH!! would you please just answer one more thing?

OpenStudy (johnweldon1993):

of course :)

OpenStudy (anonymous):

If I have to find the mode of decimal numbers, how can I find it if no numbers are repeated?

OpenStudy (johnweldon1993):

Then there would be no mode :)

OpenStudy (anonymous):

12.5,12.7,14.9,11.3,11.7,10.9, is there no mode here? :)

OpenStudy (johnweldon1993):

There is none indeed :)

OpenStudy (anonymous):

THANKYOU SO MUCH JOHN :) I really appreciate it!

OpenStudy (johnweldon1993):

Anytime hun :)